OpenStudy (anonymous):

im stuck on h7p2 could someone help with a little guidance?

OpenStudy (anonymous):

Part a) Short circuit L and evaluate Vo as a voltage divider b)Time constant is equal to T=L/(Rnorton) c)use equations V0=I(t)*R0 and I(t)-V/(Ro+RS)*(1-e^((t*Ro+RS)/L) to solve for t d) they give you the speed of light 299792458m/s. and 2cm so solve (distance/time = speed of light) Hope that helps

OpenStudy (anonymous):

thanks cot, it really helped me on d, on a,b,c im still triying to figure them out.

OpenStudy (anonymous):

Hey Cotdamn, thank you very much i know finished all h7p1 and h7p2 :)

OpenStudy (anonymous):

Sweet! glad to hear that you got the green check mark